OpenStudy (anonymous):

In an autocatalytic chemical reaction, the product formed is a catalyst for the reaction. If Qo is the amount of the original substance and x is the amount of catalyst formed, the rate of chemical reaction is Q'(x)= kx(Q0 − x) For what value of x will the rate of chemical reaction be greatest?

OpenStudy (anonymous):

ok, so I have no idea how to deal with this. i should find where Q''(x)=0, right? but how do you take the derivative with all those variables (still on my 1st semester).

OpenStudy (inkyvoyd):

Those variables that are constant are treated as constants. Those variables that are variables are treated as functions of whatever you are taking the derivative to the respect to. Use the chain rule. I'm pretty sure only the first sentence applies to you, however.

OpenStudy (anonymous):

so is the derivative -2kx+Qo

OpenStudy (inkyvoyd):

I think so.

OpenStudy (inkyvoyd):

I've never taken a formal calculus course, so you probably should wait for another to answer, like @satellite73

OpenStudy (anonymous):

or @UnkleRhaukus or @Zarkon ? any of you out there?

OpenStudy (anonymous):

i know nothing about this, but you are given the derivative, it is \[Q'(x)= kx(Q_0 − x)\]

OpenStudy (anonymous):

you want the max, set it equal to zero and solve

OpenStudy (anonymous):

ok that is wrong, you want the max of the rate take the derivative of this mess, set it equal to zero and solve

OpenStudy (anonymous):

you have \[Q'(x)=Q_0kx-kx^2\] max is at the vertex, namely \[-\frac{b}{2a}=-\frac{Q_0k}{2\times -k}=\frac{A_0}{2}\]

OpenStudy (anonymous):

\[-\frac{b}{2a}=-\frac{Q_0k}{2\times -k}=\frac{Q_0}{2}\]

OpenStudy (anonymous):

my typing is not so good

OpenStudy (anonymous):

ok, that makes more sense. I'm with ya there

OpenStudy (inkyvoyd):

satellite, they are asking for the rate of the rate, which is a double derivative

OpenStudy (anonymous):

so would I solve for Qo in 0=kx-kx^2 and then put it into Qo/2 to get the max?

OpenStudy (anonymous):

Sorry, 0=Qo(kx)-kx^2

OpenStudy (inkyvoyd):

hick, they want when the rate is the greatest.

OpenStudy (inkyvoyd):

That means that the rate of the rate must be 0

OpenStudy (inkyvoyd):

0=-2kx+Qo

OpenStudy (anonymous):

ok, so x=Qo/2k

OpenStudy (inkyvoyd):

yup.

OpenStudy (anonymous):

but they're asking me for an x value

OpenStudy (anonymous):

and they don't tell you Qo or k

OpenStudy (anonymous):

that is the x value

OpenStudy (anonymous):

and that is what you are asked for, so you are done

OpenStudy (anonymous):

my homework won't accept that answer

OpenStudy (anonymous):

am I missing something important here? I gave you everything they gave me for the question. I wonder what else they could be looking for?

OpenStudy (unklerhaukus):

The rate of chemical reaction will be greatest when \[x=\frac {Q_0}{2}\]

OpenStudy (anonymous):

yay! it took that! i wonder where the k went?

OpenStudy (anonymous):

Thanks a bunch!

OpenStudy (unklerhaukus):

I just read and re-read the post above made by satellite73, i dont think they are asking for the rate of the rate